A 0.045 kg aluminum bullet traveling at 471 m/s strikes an armor plate and comes to a stop. If all its energy is converted to heat that is absorbed by the bullet, what is the bullet's temperature change in degrees Celsius?
Answer #1

Q absorbed = KE


Q = m*C*dT = 0.5*m*v^2

dT = (0.5*v^2)/(C)

dT = (0.5*471*471)/870 = 127.49 oC